Given g(x) = 5x – 4, write the value of g(2a)
Found 2 solutions by junior403, Nate:
Answer by junior403(76)

Given g(x) = 5x – 4, write the value of g(2a)
This represents a function where (2a) is substituted for the value of (x).
So...
If g(x) = 5x - 4
Then...
g(2a) = 5(2a) - 4
We solve this in terms of a.
So first we multiply 5(2a)...
= 10a - 4
Subtract 4 from both sides of the equation...
4 = 10a
Then divide both sides by 10...
4/10 = a
Then simplify the expression...
2/5 = a
Good Luck!
Thanks for the question!

Answer by Nate(3500)

g(x) = 5x – 4
g(2a) = 5(2a) – 4
g(2a) = 10a - 4
There is no value for 'a' because we do not know the value for g(2a). So, the answer is 10a+-+4